1)a solid rectangular block of copper 5cm by 4cm by 2cm is drawn out to make a cylindrical wire of diameter 2 mm. Calculate the length of wire.
2)water flows through a circular pipe of internal diameter 3cm at a speed of 10cm/s. If the pipe is full, how much water flows from one pipe in one minute ( answer in liters)
Show working out.

Answer by mananth(16949)

Volume of copper = 4*5*2= 40cm^3
...
radius of wire =0.1 cm
..
Volume of wire = pi*r^2*l
40 = pi*0.1^2*l
40/(pi*0.1^2)= l
length = 1273 cm
...
volume of pipe = pi*r^2*l
V= pi*1.5^2*l0
70.68 cm^3 /second
..
1cm^3= 0.001 liter
70.68 cm^3= 0.07068 liters
..
This flow is per second
for 60 second it will be 60 * 0.07068
..
the flow will be 4.24 liters /minute
m.ananth@hotmail.ca